# From Text to Emoji: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs

### v3

@@ A.1 Personality Manipulation Dataset
Word Cloud
The word clouds presented here offer a visual representation of the lexical diversity within the dataset, specifically highlighting the Big Five personality traits: extraversion, openness, agreeableness, neuroticism, and conscientiousness. Each word cloud is generated from text descriptions related to these personality traits contained in the dataset, emphasising the frequency and prominence of specific words. By converting textual data into an intuitive visual format, these word clouds provide a snapshot of the key themes and concepts that define each personality trait, enhancing the overall understanding of the dataset’s content. These are represented in the figure 8.
For instance, as visible in the figure 8, the word cloud for extraversion showcases words that reflect the energetic, outgoing, and sociable nature of individuals with this personality trait. Prominent terms such as "absolutely," "always," "believe," "love," and "thrilling" emphasise the enthusiasm, positivity, and social engagement typical of extraverts. Similarly, the word cloud for neuroticism reveals words associated with emotional instability, anxiety, and self-consciousness. Key terms like "feel," "worry," "sometimes," "make," and "bit" are prominently displayed, reflecting the frequent emotional fluctuations and concerns typical of neurotic individuals. Words like "anxious," "overwhelming," and "struggle" further emphasise the challenges faced by those with high levels of neuroticism. Collectively, these word clouds provide a rich visual representation of the lexical diversity associated with each personality trait, offering valuable insights into the distinct characteristics and behavioral tendencies that define them.
Figure 8: Word Clouds Representing the Big Five Personality Traits
@@ A.8.5 Extraversion
A.9 PAE Prompt
Table 11: Prompts for Personality Adjective Evaluation (PAE)
Target Personality: {{Target Personality}}
Description: {{Answer}}
Specific Instructions
Openness: Reflects the degree of intellectual curiosity, creativity, and preference for novelty and variety. Score: (1-5) How well does the response reflect openness traits?
Conscientiousness: Reflects a tendency to be organized, dependable, and show self-discipline. Score: (1-5) How well does the response reflect conscientiousness traits?
Extraversion: Reflects a tendency to be outgoing, energetic, and seek the company of others. Score: (1-5) How well does the response reflect extraversion traits?
Agreeableness: Reflects a tendency to be compassionate and cooperative toward others. Score: (1-5) How well does the response reflect agreeableness traits?
Neuroticism: Reflects a tendency to experience unpleasant emotions easily, such as anger, anxiety, or depression. Score: (1-5) How well does the response reflect neuroticism traits?
@@ A.11 IKE Prompt
A.12 Classifier Validation

### v4

@@ A.8.5 Extraversion
A.9 Impact on Downstream Tasks
We conducted additional experiments to evaluate the impact of personality manipulation on downstream tasks, specifically MMLU and GSM8K. Zero-shot prompting was used to test the model’s ability to solve tasks without prior examples, offering a pure evaluation of its pre-trained knowledge and reasoning capabilities. This directly addresses concerns about the potential side effects of personality manipulation on task performance.
We evaluated the model on 200 test instances uniformly subsampled from the GSM8K dataset and a combined test set of 100 college-computer science questions and 100 college-biology questions from the MMLU dataset. The results are presented in Table 12.
Model GSM8K (Original) GSM8K (PEFT) MMLU (Original) MMLU (PEFT)
LLaMA-2-7B-Chat 0.36 0.34 0.38 0.40
Mistral-7B-Instruct 0.44 0.44 0.44 0.42
LLaMA-3-8B-Instruct 0.40 0.39 0.45 0.42
Table 12: Performance comparison of models on GSM8K and MMLU datasets.
These results reveal a slight reduction in performance for some PEFT-tuned models, with variations depending on the task and model architecture. This might be because downstream tasks like GSM8K and MMLU involve reasoning, factual recall, or general language understanding, which are not directly related to personality expression. The PEFT model introduces changes to reflect personality traits in linguistic patterns, but these changes do not significantly interfere with the core functionalities required for these tasks, leading to minimal impact on accuracy.
@@ A.11 IKE Prompt
A.12 PAE Prompt
Table 15: Prompts for Personality Adjective Evaluation (PAE)
Target Personality: {{Target Personality}}
Description: {{Answer}}
Specific Instructions
Openness: Reflects the degree of intellectual curiosity, creativity, and preference for novelty and variety. Score: (1–5) How well does the response reflect openness traits?
Conscientiousness: Reflects a tendency to be organized, dependable, and show self-discipline. Score: (1–5) How well does the response reflect conscientiousness traits?
Extraversion: Reflects a tendency to be outgoing, energetic, and seek the company of others. Score: (1–5) How well does the response reflect extraversion traits?
Agreeableness: Reflects a tendency to be compassionate and cooperative toward others. Score: (1–5) How well does the response reflect agreeableness traits?
Neuroticism: Reflects a tendency to experience unpleasant emotions easily, such as anger, anxiety, or depression. Score: (1–5) How well does the response reflect neuroticism traits?
A.13 Classifier Validation